the dosage of technetium-99m for a lung scan is 20microCi/kg of body mass. how many milicuries of technetium-99m should be given to a 48 kg person?


media%2F384%2F38406283-a6f6-4d8e-88cc-15
Answer #1

1 micro Ci = 0.001 mili Curie

thus

20 micro Ci = 20 * 0.001 = 0.02 mili Ci

dosage of technetium-99m for a lung scan is 20 microCi / kg of body mass = 0.02 mili Ci / Kg body mass.

we have

mass of person = 48 g

thus

milicuries of technetium-99m should be given = 0.02 * 48 = 0.96 mili Ci (answer)